The area of a square is between 16 and 60 inclusively. What are all the possible integer values for the length of the sides of the square?
Answer by Theo(13342)

16 < s^2 < 60
Let s = length of one of the sides of your square.
If s = 4, then s^2 = 16
If s = 5, then s^2 = 25
If s = 6, then s^2 = 36
If s = 7, then s^2 = 49
If s = 8, then s^2 = 64

Looks like your integers would be:

4,5,6,7